OpenStudy (anonymous):

write in summation notation 125/9+625/16+3125/25+15,625/36+...

OpenStudy (anonymous):

"Summary Notation"

OpenStudy (anonymous):

Divorce the numerators from the denominators, you'll find that the numerators increase by multiplying 5 over and over again, starting at 125, or 5 cubed :D The denominators are just the squares of all the numbers starting at 3. :)

OpenStudy (anonymous):

This would seem about right... \[\huge \sum_{n=3}^{\infty}\frac{5^n}{n^2}\]

OpenStudy (anonymous):

what if n = 1 ? cause the question says n should equal 1

OpenStudy (anonymous):

You want it to start at n=1? :D That can be arranged... \[\huge \sum_{n=1}^{\infty}\frac{5^{n+2}}{(n+2)^2}\]

OpenStudy (anonymous):

Thank you very much